Shelias measured glucose level one hour after a sugary drink varies according to the normal distribution with u 124 mg/dl and σ-10.2 mg/dl. What is the level L such that there is probability only 0.01 that the mean glucose level of 4 test results falls above L? Give your answer precise to one decimal place
Answer #1

Solution :  

\mu\bar x = 124

\sigma\bar x = \sigma / \sqrt n = 10.2 / \sqrt 4 = 5.1

P(Z > 2.326) = 0.01

z = 2.326

\bar x = z * \sigma \bar x+ \mu \bar x = 2.326 * 5.1 + 124 = 135.9

L = 135.9